Use the method of substitution to solve the system.
x^2+y^2=8
x+y=0

Answer by unlockmath(1688)

Hello,
With this we can do the following:
x^2+y^2=8
x+y=0
The second equation subtract y from both sides to get:
x=-y
Now substitute this into the first equation:
(-y)^2+y^2=8
This can be rewritten as:
2y^2=8
Now divide by 2 to get:
y^2=4
Square root both sides to get:
y=+-2
x=+-2
Make sense?
RJ
www.math-unlock.com